# Constraints to Lorentz violation and ultrahigh-energy electrons in D-foamy space-times

### v1

@@ Acknowledgements.
C.L. is indebted to an anonymous reviewer of his Ph.D. Thesis for useful remarks, regarding foam effect on charged leptons, that have partially inspired the work presented in this paper. This work is supported by the National Natural Science Foundation of China under Grants No. 12335006 and No. 12075003. The work of C.L. is partly supported by a Boya Fellowship provided by Peking University, and by the China Postdoctoral Science Foundation (CPSF) under Grant No. 2024M750046. C.L. gratefully acknowledges the financial support received from Postdoctoral Fellowship Program (Grade B) of CPSF under Grant No. GZB20230032. The work of C.L. also falls within the scope of the COST Action CA23130 -“Bridging high and low energies in search of quantum gravity” (BridgeQG).
@@ Appendix A Implication of brane recoil on space-time
Consider a D-particle (0-brane), initially at rest at $Y^{i}(t=0)\equiv Y_{i}$ of spatial coordinates of a $(D+1)$-dimensional space-time (which could be a 3-brane world), experiencing at $t=0$ an impulse due to scattering off a matter string. For heavy defects Ellis:1999uh its trajectory, $Y^{i}(t)$, is specified by (56), where the vertex operator, corresponding to the present case with the relevant background $g^{J}=u_{i}$, is
@@
Eq. (57) contains actually a pair of deformations, which are relevant in a worldsheet renormalization group (RG) sense, having anomalous dimensions $-\varepsilon^{2}/2$, where $\varepsilon$ is a regulator. They form a modified (logarithmic) conformal algebra (superconformal algebra, in the case of superstrings), which closes if and only if, one links Mavromatos:1998nz $\varepsilon^{-2}$ with the ‘running’ RG scale, $\ln A$, given in terms of the worldsheet area $A$:

### v2

@@ Acknowledgements.
C.L. is indebted to an anonymous reviewer of his Ph.D. Thesis for useful remarks, regarding foam effect on charged leptons, that have partially inspired the work presented in this paper. This work is supported by the National Natural Science Foundation of China under Grants No. 12335006 and No. 12075003. The work of C.L. is partly supported by a Boya Fellowship provided by Peking University, and by the China Postdoctoral Science Foundation (CPSF) under Grant No. 2024M750046. C.L. gratefully acknowledges the financial support received from Postdoctoral Fellowship Program (Grade B) of CPSF under Grant No. GZB20230032. The work of C.L. also falls within the scope of the COST Action CA23130 – Bridging high and low energies in search of quantum gravity (BridgeQG).
@@ Appendix A Implication of brane recoil on space-time
Consider a D-particle (0-brane), initially at rest at $Y^{i}(t=0)\equiv Y_{i}$ of spatial coordinates of a $(D+1)$-dimensional space-time (which could be a 3-brane-world), experiencing at $t=0$ an impulse due to scattering off a matter string. For a heavy defect Ellis:1999uh its trajectory, $Y^{i}(t)$, is specified by (56), where the vertex operator, corresponding to the present case with the relevant background $g^{J}=u_{i}$, is,
@@
Eq. (57) contains actually a pair of deformations, which are relevant in a worldsheet renormalization group (RG) sense, having anomalous dimensions $\varepsilon^{2}/2$, where $\varepsilon$ is a regulator. They form a modified (logarithmic) conformal algebra (superconformal algebra, in the case of superstrings), which closes if and only if, one links Mavromatos:1998nz $\varepsilon^{-2}$ with the ‘running’ RG scale, $\ln A$, given in terms of the worldsheet area $A$: